OpenStudy (anonymous):

HI everyone! Can someone help me solve the inverse laplace transform of -(1/4)L^-1{1/(s+2)^3) please? I know the answer is -(1/8)t^2e^-2t but I need help getting there. Thanks! :o)

OpenStudy (anonymous):

I know 1/(s+2)^3 resembles n! / s^(n+1) so n would be equal to 2... then you have to multiply the top and bottom by a fancy form of 1, in the case, 2factorial/2factorial (2!/2!)... so then the problem would look like this... -(1/4)(1/2!) L^-1 { 2/(s+2)^(2+1)}... THEN...

OpenStudy (anonymous):

i also know that L^-1{n!/s^(n+1)} = t^n this lead me to want to make the answer... -(1/4)(1/2!) times t^2 which would make the answer -(1/8)t^2 but that's not the answer...so I need help with this step how do I get from -(1/4)(1/2!) L^-1 { 2/(s+2)^(2+1)} to -(1/8)t^2e^-2t ???

OpenStudy (anonymous):

@rational @Michele_Laino if anyone can help, please do

OpenStudy (michele_laino):

do you want to know how to get the inverse transform of this function: \[\Large \frac{1}{{{{\left( {s + 2} \right)}^3}}}\]

OpenStudy (anonymous):

yes basically...not forgetting the -1/4 in front...but yes

OpenStudy (anonymous):

i typed everything above pretty neatly...you can see the step I get stuck on if you take a second and review my work above :o)

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

I also tried to explain my thought process as I was typing as well

OpenStudy (michele_laino):

In general the subsequent identity holds: \[\Large L\left( {{t^n}{e^{ct}}} \right) = \frac{{n!}}{{{{\left( {s - c} \right)}^{n + 1}}}}\]

OpenStudy (anonymous):

one mistake I made... this L^-1{n!/s^(n+1)} = t^n should have been L^-1{t^n}=n!/s^(n+1)

OpenStudy (anonymous):

michele...where did you get that transform? from some table?

OpenStudy (michele_laino):

no, we can demonstrate it, using the concept of differetiation of the L-transform

OpenStudy (anonymous):

uhm

OpenStudy (michele_laino):

for example we have to start from the definition of L-transform, like this: \[L\left( {f\left( t \right)} \right) = \int_0^\infty {f\left( t \right){e^{ - st}}dt} \]

OpenStudy (michele_laino):

\[\Large L\left( {f\left( t \right)} \right) = \int_0^\infty {f\left( t \right){e^{ - st}}dt} \]

OpenStudy (anonymous):

okay...you don't have to go through that integral...i get it...it will equal t^2e^at

OpenStudy (anonymous):

but my problem is this...

OpenStudy (anonymous):

when you are working on an inverse laplace transform, you most likely will not have that integral and its answer sitting around...I need to know how to work backwards towards that without actually knowing it

OpenStudy (michele_laino):

please it is my starting point. That function is a function of the variable s, so we can rewrite it as below: \[\Large F\left( s \right) = \int_0^\infty {f\left( t \right){e^{ - st}}dt} \] Now we have to consider, the subsequent incremental ration: \[\Large \frac{{F\left( {s + \sigma } \right) - F\left( s \right)}}{\sigma } = \int_0^\infty {\frac{{{e^{ - \sigma t}} - 1}}{\sigma }{e^{ - st}}f\left( t \right)dt} \]

OpenStudy (anonymous):

how do I get from -(1/4)(1/2!) L^-1 { 2/(s+2)^(2+1)} to -(1/8)t^2e^-2t ??? is it possible or do you ABSOLUTELY HAVE TO KNOW that integral?

OpenStudy (michele_laino):

ratio*

OpenStudy (anonymous):

i don't even know what a subsequent incremental ratio is...i never studied that

OpenStudy (michele_laino):

ok! the general formulas are these: the inverse transform of this function: \[\Large \frac{1}{{{{\left( {s - a} \right)}^k}}}\] is: \[\Large \frac{1}{{\left( {k - 1} \right)!}}{t^{k - 1}}{e^{at}}\]

OpenStudy (anonymous):

my teacher only required us to remember 7 basic transforms...i have never seen this before

OpenStudy (michele_laino):

we can demonstrate that formula recursively

OpenStudy (anonymous):

i don't even know what that means lol

OpenStudy (anonymous):

what is recursively?

OpenStudy (michele_laino):

it is simple, since those formulas depend on natural parameter k, then we can apply the mathematical induction principle to k

OpenStudy (anonymous):

i don't even know what the mathematical induction principal is michele i don't think you are going to be able to help me...too much "mathese" to understand

OpenStudy (anonymous):

here is what i know...the top 8 transforms plus the hyperbolics on this page...that's it http://www.egr.msu.edu/~priezjev/ME391/Laplace.pdf

OpenStudy (michele_laino):

you can memorize those formulas, then

OpenStudy (anonymous):

i have them memorized...i just need to know how to get from -(1/4)(1/2!) L^-1 { 2/(s+2)^(2+1)} to -(1/8)t^2e^-2t

OpenStudy (michele_laino):

all those formulas, in your table can be demonstrated usin the concept of derivative of L-transform, and the mathematical induction principle

OpenStudy (michele_laino):

using*

OpenStudy (anonymous):

okay but why are you telling me that? i don't understand...does that help me solve my problem somehow?

OpenStudy (michele_laino):

because in order to solve your problem I have to explain you how I'm able to get some formulas

OpenStudy (anonymous):

i don't even know what concept of derivative of L-transform is...is that having a laplace trransform sitting on your paper, then taking its derivative or something?

OpenStudy (rational):

Notice that \(\large \mathcal{L}^{-1}(\frac{1}{s^3})= \frac{t^2}{2!}\)

OpenStudy (rational):

familiar with exponential shift formula ?

OpenStudy (rational):

exponential shift formula : \[\mathcal{L}(e^{at}f(t))=F(s-a)\]

OpenStudy (anonymous):

wait...exponential shift...that is the same thing as the first translation theorem right?

OpenStudy (rational):

Yep!

OpenStudy (anonymous):

okay great but wait...

OpenStudy (rational):

k

OpenStudy (anonymous):

that 1st translation theorem is in chapter 7.3 the problem I am trying to solve is in 7.2 I shouldn't have to use that

OpenStudy (rational):

Ohk

OpenStudy (anonymous):

so frustrated with this problem...how about I take a pic of the book...it works out the problem...then maybe you can just tell me how they did it eh?

OpenStudy (rational):

exponential shift formula is right after the definition of laplace transform in my textbook

OpenStudy (rational):

sure post it

OpenStudy (anonymous):

well my book is dumb then lol

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

the top left says L^-1

OpenStudy (rational):

also attach example6 of section 7.1

OpenStudy (anonymous):

it's just a long old integral

OpenStudy (rational):

then they're just using the previous result thats all there is to it.

OpenStudy (anonymous):

so they just magically expect me to recall an obscure integral from a previous section and "know" to apply it?

OpenStudy (rational):

For now yes once you know "exponential shift formula", you don't need to memorize these it seems they're covering it in next section itself, so have patience

OpenStudy (anonymous):

omg! gurrrrrr

OpenStudy (anonymous):

so are you telling me without knowing the answer to that integral, there is no "mechanical" way to solve this problem?

OpenStudy (rational):

But wait a second, you're actually using "exponential shift formula" for calculating the inverse of 1/(s+2) right ?

OpenStudy (anonymous):

i don;'t know...i would have to think about that for a sec

OpenStudy (rational):

how do you know \(\mathcal{L}^{-1}(\frac{1}{s+2})=e^{-2t}\) ?

OpenStudy (anonymous):

because L{e^at} = 1/s-a

OpenStudy (anonymous):

so a=-2

OpenStudy (rational):

good, then there is a nice trick to work \(\mathcal{L}^{-1}(\frac{1}{(s+2)^3})\)

OpenStudy (rational):

lets see

OpenStudy (rational):

forget about inverses for sometime

OpenStudy (anonymous):

from that I know n=2 and I also know you have to multiply by a fancy form of 1, in this case 2factorial/2factorial

OpenStudy (anonymous):

so I would have -(1/4)(1/2!) L^-1 { 2/(s+2)^3}

OpenStudy (rational):

we know that \[\mathcal{L}(1)=\frac{1}{s}\] then consider \(\mathcal{L}(e^{at}f(t))\)

OpenStudy (rational):

\(f(t)\) is just some arbitrary function

OpenStudy (anonymous):

yes

OpenStudy (rational):

\[\mathcal{L}(e^{at}f(t))=\int\limits_0^\infty e^{at}f(t)e^{-st}\,dt\] yes ?

OpenStudy (anonymous):

yes

OpenStudy (rational):

\[\begin{align}\mathcal{L}(e^{at}f(t))&=\int\limits_0^\infty e^{at}f(t)e^{-st}\,dt\\~\\ &=\int\limits_0^\infty f(t)e^{-(s-a)t}\,dt\\~\\ &=F(s-a) \end{align}\] yes ?

OpenStudy (anonymous):

yes...I've actually worked that integral 3 times in the past...so yes

OpenStudy (rational):

Sweet :) we're almost done then

OpenStudy (rational):

we know that \(\large \mathcal{L}(t^2)=\frac{2!}{s^3}\) can you guess the value of \(\large \mathcal{L}(e^{at}t^2)\) ?

OpenStudy (anonymous):

do you mean the inverse laplace?

OpenStudy (rational):

no, laplace transform only

OpenStudy (rational):

use the previous result http://gyazo.com/8ebe92d7699329f175da0ef7da0f1014

OpenStudy (anonymous):

nevermind what I said...getting inverses mixed up

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

uhmmm

OpenStudy (rational):

we know that \(\large \mathcal{L}(t^2)=\frac{2!}{s^3}\) and we also know that \(\large \mathcal{L}(e^{at}f(t)) = F(s-a)\) therefore \(\large \mathcal{L}(e^{at}t^2) = \frac{2!}{(s-a)^3}\)

OpenStudy (rational):

basically multiplying any function by\(e^{at}\) shifts its transform by \(a\)

OpenStudy (anonymous):

okay, but without knowing the 1st translation theorem, I still can't mechanically solve this problem

OpenStudy (rational):

when you don't know anything, it is good because you don't have a choice... u go and use the definition of laplace transform

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay my turn...

OpenStudy (anonymous):

look at the right hand side of that problem I posted...its an inverse correct?

OpenStudy (anonymous):

without knowing the 1st translation theorem and without mysteriously remembering that crazy integral...can you solve the problem I posted?

OpenStudy (anonymous):

can you mechanically solve it just knowing the basic transforms?

OpenStudy (rational):

Ahh nope, it wont reduce further in partial fraction decomposition if you're not allowed to use 1st translation rule/convolution stuff, you just need to evaluate the laplace transform integral of \(e^{at}t^2\) and remember it somehow

OpenStudy (rational):

but you're thinking too much on this small problem memorizing these is unnecessary once you get to know the 1st translation rule

OpenStudy (anonymous):

yeah but what I am saying is that if you only have the problem to look at, how on earth would you know to apply the e^att^2 thing...that seems almost impossible of a leap